OpenStudy (anonymous):

Choose the correct simplification of the expression (2b/c)^3

OpenStudy (anonymous):

\[(2b/c)^{3}=(2^{3}b ^{3}/c ^{3})= 8b ^{3}/c ^{3}\] Is that one of the choices?

OpenStudy (zzr0ck3r):

(2b)^3/c^3 (2^3*b^3)/c^3 = (8*b^3/c^3

OpenStudy (anonymous):

Thank you took me a while but I actually ended up solving it.